OpenStudy (anonymous):

What is the slope of the line that passes through the points (-1, -17) and (-4, -2)?

OpenStudy (anonymous):

these are the options Option A: 5 Option B: 1/5 Option C: -1/5 Option D: -5

OpenStudy (anonymous):

Haha owk (: give me a second

OpenStudy (anonymous):

(-17 + 2)/(-1 +4) = -15/3 = -5

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

The equation is deltaY/deltax i.e (y1 - y2)/(x1-x2)